OpenStudy (anonymous):

For the function f(x) = 3x+2, find x such that f(x) = 14. So with this equation, do they just mean to fill in 14 for x?

jimthompson5910 (jim_thompson5910):

No, they want you to replace f(x) with 14 and solve for x

OpenStudy (precal):



OpenStudy (anonymous):

How do I solve for x? @precal and @jim_thompson5910

jimthompson5910 (jim_thompson5910):

14 = 3x + 2 14-2 = 3x 12 = 3x 3x = 12 x = 12/3 x = 4

jimthompson5910 (jim_thompson5910):

The idea is to basically follow the order of operations in reverse to isolate x

OpenStudy (anonymous):

I got it! Thanks!

jimthompson5910 (jim_thompson5910):

sure thing